# Two-component relativistic equation-of-motion coupled cluster for electron ionization

### v1

@@ V Conclusions
We have implemented relativistic formulations of IP-EOMCCSD, IP-EOMCCSD(3h2p), and IP-EOMCCSDT where the relativistic treatment is based on the 1eX2C or DCB-X2C formalisms. A basis set study on vertical IPs for HCCX [X = Cl, Br, I] molecules at the IP-EOMCCSD level revealed surprising discrepancies between results obtained from the 1eX2C- or DCB-X2C-based calculations when using small contracted basis sets. At the double-$\zeta$ level, 1eX2C- and DCB-X2C-IP-EOMCCSD calculations carried out using contracted x2c- or ANO-RCC-type basis sets produce IP values that differ by up to $\sim$0.1 eV. Calculations using uncontracted Dyall-, x2c-, and ANO-RCC-type basis sets eliminate these differences, which suggests that the error stems from the recontraction of the x2c- and ANO-RCC-style basis sets at the 4c to 2c transformation part of the algorithm. These recontraction issues appear to be less of a problem at larger $\zeta$-levels.
We have also shown that the inclusion of correlation effects beyond those captured by IP-EOMCCSD are essential for achieving accurate estimates of IP values in the HCCX series. A DCB-X2C-based composite approach that corrects adiabatic IP-EOMCCSD energetics obtained with a quadruple-$\zeta$ basis set with IP-EOMCCSDT data obtained in a double-$\zeta$ basis set achieves errors relative to experiment on the order of 0.1 eV or less. Remaining error in the computed adiabatic IP values could be attributed to the use of inexact geometries for the neutral and cation species or to the modest size of the basis set used for the IP-EOMCCSDT correction. Large-basis X2C-IP-EOMCCSDT calculations may be challenging to carry out in general, in which case approaches such as the active-space IP-EOMCCSDt approach,[41, 42] could be a viable alternative to the composite scheme we have employed.
Supporting Information Comparison between IP values obtained from all-electron and frozen-core IP-EOMCCSD/x2c-TZVPPall-2c calculations.

### v2

@@ V Conclusions
[…]
We have also shown that the inclusion of correlation effects beyond those captured by IP-EOMCCSD are essential for achieving accurate estimates of IP values in the HCCX series. A DCB-X2C-based composite approach that corrects adiabatic IP-EOMCCSD energetics obtained with a quadruple-$\zeta$ basis set with IP-EOMCCSDT data obtained in a double-$\zeta$ basis set achieves errors relative to experiment on the order of 0.1 eV or less. Remaining error in the computed adiabatic IP values could be attributed to the use of inexact geometries for the neutral and cation species or to the modest size of the basis set used for the IP-EOMCCSDT correction. Large-basis X2C-IP-EOMCCSDT calculations may be challenging to carry out in general, in which case approaches such as the active-space IP-EOMCCSDt approach,[41, 42, 43] could be a viable alternative to the composite scheme we have employed.
Our calculations also demonstrate the effectiveness of the 1eX2C relativistic framework combined with DCB-parameterized SNSO scaling factor in replicating results from the more rigorous DCB-X2C approach; the IP values obtained by both methods are quite similar, especially in the limit of large basis sets. Furthermore, 1eX2C does provide non-negligible savings in memory usage and computational time. For example, in the IP-EOMCCSD calculation for HCCI with Dyall v4z basis set, 1eX2C-IP-EOMCCSD offers a reduction of 21% in total wall time compared to DCB-X2C-IP-EOMCCSD (see Table S3 in the Supplemental Information), while producing essentially identical IP values and $\Delta_{\Omega}$ splittings. The computational time savings become 37%–38% when the calculations for HCCCl and HCCBr with cc-pV5Z-DK are considered, due to the use of recontracted basis set in the SCF step. Of course, these savings will lessen when considering higher-order correlation effects. Nevertheless, the 1eX2C+SNSO approach appears to be a reasonable alternative to DCB-X2C, particularly when considering larger molecular systems and contracted basis sets.
[…]
@@ Acknowledgements.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.